OpenStudy (anonymous):

8+3n over 12=13? how do i show my work?

OpenStudy (anonymous):

Just write what you did.\[\frac{ 8 + 3n }{ 12 } = 13\] Multiply both sides by 12.\[\frac{ 8 + 3n }{ 12 } (12)= 13(12)\] 8 + 3n = 156 Subtract 8 from both sides. 3n = 356 Divide by 3. n = 356 /3 Solve for n.

OpenStudy (anonymous):

@simplya